OpenStudy (jack1):

Would really love some help understanding where to start please... : Using eigenvalues and eigenvectors, find the general solution of the following system of equations: dx/dt=-4x+2y dy/dt=-5/2 x+2y

OpenStudy (jack1):

sorry, this (dy/dt=-5/2 x+2y) should look like this: \[\frac{ dy }{ dt } = -\frac{ 5 }{ 2 }x +2y\]

OpenStudy (perl):

thats find

OpenStudy (perl):

fine**

OpenStudy (perl):

i have to read up on this http://tutorial.math.lamar.edu/Classes/DE/LA_Systems.aspx

OpenStudy (perl):

if you click it , he has stuff on eigenvectors relating to DE equations

OpenStudy (jack1):

sorry, was just finding what i could on wolfram, back now thanks man, just headin to the link now